# Lê modules and hypersurfaces with one-dimensional singular sets

### v2

@@ Abstract.
By using our previous results on Lê modules and an upper-bound on the betti numbers which we proved with Lê, we investigate the cohomology of Milnor fibers and the internal local systems given by the vanishing cycles of hypersurfaces with one-dimensional singular sets.
@@ 1. Introduction
Furthermore, ${\operatorname{mult}}_{\mathbf{0}}f=2$ and $m=2$, and if $r=1$, then $\lambda^{1}=2$, and so $\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}$.
@@ 5. Remarks
Throughout this article, we have allowed for torsion in the cohomology of the Milnor fiber. In fact, we do not know if such torsion can possibly exist when $\dim_{\mathbf{0}}\Sigma f=1$.
However, when $\dim_{\mathbf{0}}\Sigma f=2$, there are the famous examples of Cohen, Denham, and Suciu in [3] of hypersurfaces with torsion in the cohomology of the Milnor fiber.

### v3

@@ Abstract.
By using our previous results on Lê modules and an upper-bound on the betti numbers which we proved with Lê, we investigate the cohomology of Milnor fibers and the internal local systems given by the vanishing cycles of hypersurfaces with one-dimensional singular sets and small Lê numbers.
@@ 1. Introduction
Furthermore, if $r=1$, then $\lambda^{1}=2$, and so $\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}$.
@@ 5. Remarks
Throughout this article, we have allowed for torsion in the cohomology of the Milnor fiber. One might wonder if this can actually occur when $\dim_{\mathbf{0}}\Sigma f=1$.
In fact there is a relatively recent example of such a phenomenon given in [15], which is discussed in a very illuminating way by Suciu in 11.3 of [14]. This example consists of 16 (reduced) hyperplanes through the origin in ${\mathbb{C}}^{3}$, and the critical locus consists of 45 lines. It is shown that there is $2$-torsion in the degree 1 homology group of the Milnor fiber, i.e., in the degree 2 cohomology group. Our results in this paper for functions with low Lê numbers say essentially nothing about this example (though one can use the formulas in Proposition 1.4 of [10] to show that $\lambda^{1}=165$ and $\lambda^{0}=900$).
However we do not know if such torsion can possibly exist when $\Sigma f$ is a smooth curve or even has a single irreducible component.